OpenStudy (anonymous):

Be sure to simplify by collecting like terms before you solve. 8x - 5x = -15

OpenStudy (anonymous):

what do you think ur like terms are?

OpenStudy (anonymous):

8x and -5x

OpenStudy (anonymous):

correct! so combine those and what do you get? 8x-5x= ???

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

 Hint: isolate x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome! :) so do u know what to do?? divide 3 from each side and you have x=-15/3.... so x= ????